Use the method of substitution to solve the system. (Enter your answer(s) as ordered pair(s). If there is more than one ordered pair, use commas between ordered pairs.)

y2−4x2=19
9y2+16x2=275

Answer by unlockmath(1688)

Hello,
With this one y2−4x2=19
9y2+16x2=275
Work with the first equation and set it up as:
y2= 4x^2 + 19
Now substitute this into the second equation:
9(4x^2+19)+16x^2=275
Expand this out:
36x^2+171+16x^2=275
Combine like terms and subtract 171 to get:
52x^2=104
Divide by 52 to get:
x^2=2
Square root both sides:
x= +-sq rt 2
Plug this into the original equation to get:
y = +-3 sq rt 3
Make sense?
RJ
www.math-unlock.com